# A YTD correspondence for constant scalar curvature metrics

### v1

@@ Abstract
Given a compact Kähler manifold, we provide a uniform YTD correspondence that characterizes the existence of a unique constant scalar curvature metric using ample/Kähler test configurations. The main point of our approach is to replace the $K$ energy with much better behaving variants, whose properties are similar to the Ding energy from the Fano case. Connections with the non-Archimedean point of view on K-stability are also briefly discussed.
@@ 1 The main results
Let $(X,\omega)$ be a compact Kähler manifold. In the polarized setting, classical formulations of the Yau–Tian–Donaldson (YTD) conjecture aim to characterize the existence of constant scalar curvature Kähler (cscK) metrics using invariants associated to test configurations [93, 85, 51]. Our work fits within this general framework. While we do not address any specific version of the YTD conjecture from the literature, we do establish a YTD correspondence: we characterize existence of cscK metrics, using ample test configurations and a novel algebraic invariant. Our approach works equally well in the transcendental setting.
@@
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X)\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
@@
where ${Ent}(u)=\frac{1}{V}\int_{X}\log(\omega^{n}_{u}/\omega^{n})\omega^{n}_{u}$ is the entropy of $\omega^{n}_{u}$ with respect to $\omega^{n}$, whereas ${I},{I}_{\textup{Ric}(\omega)}:\mathcal{H}_{\omega}\to\mathbb{R}$ are the Monge-Ampère (Aubin-Yau) energy and its Ricci-twisted version:
@@
$K^{\beta}$ energy and transcendental quantization.
@@
Recalling the formula $Ent(u)=\sup_{w\in C^{0}(X)}\left(-\log\frac{1}{V}\int_{X}e^{w}{\omega^{n}}+\frac{1}{V}\int_{X}w\omega_{u}^{n}\right)$ ([8, Proposition 2.10]) and the basic inequality $I(v)-I(u)\leq\frac{1}{V}\int_{X}(v-u)\omega_{u}^{n}$, we immediately see that $Ent^{\beta}(u)\leq Ent(u)$. In Corollary 3.11 we will see that $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
It may seem surprising at first, but our formulation of the $\beta$-entropy is intimately tied to Berman’s speculations from [6, §3.1], where he suggests to view solutions $u^{\beta}$ to the following Aubin–Yau type Monge–Ampère equation as ‘transcendental quantizations’ of $u\in\mathcal{H}_{\omega}$, with $\beta^{-1}$ serving as the Planck constant:
$(\omega+\mathrm{dd^{c}}u^{\beta})^{n}=e^{\beta(u^{\beta}-u)}\omega^{n}.$
@@
For our $\beta$-entropy, the following crucial formula is proved in Proposition 3.6:
@@
The expression on the right hand side is reminiscent of Donaldson’s $\mathcal{L}$-function from [52] in the classical Bergman kernel setting. Because of this analogy, we call $Ent^{\beta}$ the (transcendentally) quantized entropy at level $\beta$. The corresponding quantized $K^{\beta}$ energy functional $K^{\beta}:\mathcal{H}_{\omega}\to\mathbb{R}$—satisfying $K^{\beta}(u)\nearrow K(u)$ as $\beta\nearrow\infty$—is defined as follows:
@@
Due to Chen-Cheng, existence of cscK metrics is equivalent to properness of the K energy [26, 27], which we show to be further equivalent with properness of the ‘quantized’ $K^{\beta}$ functionals for $\beta$ big enough (Theorem 4.4). This is in line with quantization philosophy extensively explored in the algebraic setting (see e.g. [83, 84, 50, 52, 69, 70, 78, 96]). Our argument relies on estimates reminiscent of the partial $C^{0}$ estimate (Proposition 4.1), originally introduced by Tian [84] in the Fano setting.
In addition, we show that the $K^{\beta}$ energy and its radial version are continuous with respect to the $d_{1}$ and $d_{1}^{c}$ metrics, respectively (Lemma 3.9, Theorem 5.2). The analogous properties are known to hold for the Ding energy ([39, Proposition 5.19], [41, Theorem 1.4] [21, Proposition 3.15]) but fail for the $K$ energy. Thus, while the properness profile of $K^{\beta}$ agrees with that of the $K$ energy, its crucial continuity profile is in line with the much better behaving Ding energy known from the Fano case.
Since, in the Fano case, the favorable properties of the Ding energy have driven major advances in $K$-stability, moduli theory, $\delta$-invariants, and other areas [14, 28, 86, 60, 53, 54, 16, 91, 63, 96, 42], we are motivated to make extensive use of $K^{\beta}$ functionals in this work.
A YTD correspondence.
In the case of an ample line bundle $L\to X$ with hermitian metric $h$ whose curvature equals our Kähler metric $\omega$, to test K-stability, one considers test configurations. Historically this notion is due to Ding–Tian and Donaldson and involves one parameter degeneration of the structure $(X,L)$ [49, 85, 51]. We give the precise definition of ample test configurations in §2. One may think of these as finitely generated $\mathbb{Z}$-filtrations of the section ring $R(X,L^{k})$ for some $k\in\mathbb{N}$. This perspective was first considered in [88, 80] and it was shown in [17, Proposition 2.15] that it is equivalent with the more classical notion introduced in [85, 51].
To an ample test configuration $\mathcal{T}$, due to work initiated by Phong-Sturm, it is possible to associate a $C^{1,1}$ Mabuchi geodesic ray $0,\infty)\ni t\to u^{\mathcal{T}}_{t}\in\mathcal{H}_{\omega}^{1,1} [$72, 73, 76, 5, 30].
@@
Given that $\mathcal{T}$ is algebraic, it is natural to seek an interpretation of the associated invariants purely in algebraic terms. For instance, $J_{\mathcal{T}}$ admits such a description via certain intersection numbers attached to $\mathcal{T}$ [71, 17]. For $K^{\beta}_{\mathcal{T}}$, an explicit intersection-theoretic formula will be provided in a sequel [43] (see formula (50)). In the present work we provide a more readily available divisorial threshold–type interpretation, in the spirit of delta and alpha invariants (see Corollary 8.2), already emphasizing the algebraic nature of $K^{\beta}_{\mathcal{T}}$.
@@
To explain the proof of Theorem 1.1, let us recall a few concepts. By $(\mathcal{E}^{1},d_{1})$ we denote the $d_{1}$ metric completion of $\mathcal{H}_{\omega}$, explored in [32], with the strongly topology of the space $\mathcal{E}^{1}$ already investigated in [8]. The space $(\mathcal{E}^{1},d_{1})$ is a geodesic metric space [32, Theorem 2]. Our functionals $K$, $K^{\beta}$, and $J$ extend naturally to $\mathcal{E}^{1}$, preserving their $d_{1}$-(semi)continuity properties [7, 11]. Moreover, the extended $K^{\beta}$ continues to satisfy the formulas (4) and (6).
By $\mathcal{R}^{1}$ we denote the space of geodesic rays $t\mapsto u_{t}$ inside $\mathcal{E}^{1}$, that emanate from the background potential $0\in\mathcal{H}_{\omega}$ (notation: $\{u_{t}\}\in\mathcal{R}^{1}$). Due to Buseman convexity of $d_{1}$ ([11, Proposition 5.1]) one can introduce the chordal distance for $\{u_{t}\},\{v_{t}\}\in\mathcal{R}^{1}$ [37, 27]:
@@
Thus defined, $(\mathcal{R}^{1},d_{1}^{c})$ is a complete geodesic metric space [37, Theorem 1.4].
An important subspace of $(\mathcal{R}^{1},d_{1}^{c})$ is $\mathcal{R}^{1}_{\mathcal{I}}$, defined as the $d_{1}^{c}$-closure of rays $\{u_{t}^{\mathcal{T}}\}$ induced by test configurations $\mathcal{T}$. In [10] the authors refer to such rays simply as maximal. This space admits a natural identification with $\mathcal{E}^{1,NA}$ (see [40, Theorem 1.2] and the discussion preceding it). For more detailed treatments of this topic, we refer to [37, 10, 40, 34].
Given $\{u_{t}\}\in\mathcal{R}^{1}$ it is convenient to introduce the radial version of these functionals:
@@
In a recent exciting development, C. Li showed that rays $\{u_{t}\}\in\mathcal{R}^{1}$ satisfying $K\{u_{t}\}<\infty$ must lie in $\mathcal{R}^{1}_{\mathcal{I}}$ [61, Theorem 1.2].
The proof of Theorem 1.1 relies on four central components. From the existing literature, we will need that existence cscK metrics is equivalent with properness/geodesic stability of the K energy [26, 27], together with the aforementioned result of C. Li [61]. The remaining two components are facts we establish in this work: that the properness of the K energy is equivalent to the properness of the $K^{\beta}$ energy for sufficiently large $\beta$ (Theorem 4.4); and that the radial $K^{\beta}$ energy is $d_{1}^{c}$-continuous (Theorem 5.2).
Once the key ingredients are in place, the argument unfolds naturally. We present it here, as this simple proof serves as the skeleton of the paper (cf. [10, 53] for related ideas in the Fano case).
@@
Assume that (i) holds. By [12, Theorem 1.5] the K energy is proper, i.e., $K(u)\geq\gamma J(u)-\beta,\ u\in\mathcal{E}^{1}$ for some $\gamma,\beta>0$. By Theorem 4.4 we obtain that $K^{\beta}$ is proper too, for $\beta>0$ big enough. By our definition of $K^{\beta}_{\mathcal{T}}$ and $J_{\mathcal{T}}$, (ii) immediately follows.
Now we assume that (ii) holds. By [61, Theorem 1.2] any ray with bounded K energy slope is in $\mathcal{R}^{1}_{\mathcal{I}}$, i.e., it is approximable by test configurations. By [27] (cf. [37, Theorem 1.9]), it suffices to show existence of $\gamma>0$ such that
$\mathcal{K}\{u_{t}\}\geq\gamma J\{u_{t}\},\ \ \ \{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}.$
Let $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}$ and $\{u^{\mathcal{T}_{j}}_{t}\}$ be $d_{1}^{c}$-approximating rays of ample test configurations $\mathcal{T}_{j}$.
@@
The functional $J\{\cdot\}$ is known to be $d_{1}^{c}$-continuous (as follows from [32, Lemma 4.15] and [10, Lemma A.2]). $K^{\beta}\{\cdot\}$ also has this property by Theorem 5.2. Thus, using $K\geq K^{\beta}$ (see (19)), we can let $j\to\infty$ in the above inequality to obtain $K\{u_{t}\}\geq K^{\beta}\{u_{t}\}\geq\gamma J\{u_{t}\}.$
Given that Theorem 1.1 introduces a novel stability condition in the context of YTD type results, the second goal of the paper is to explore the meaning of $K^{\beta}$-stability, and its relationship to more classical notions of stability previously considered in the literature.
Radial $K^{\beta}$ and algebraic interpretation of $K^{\beta}_{\mathcal{T}}$.
As noted above, $K^{\beta}(u)\nearrow K(u)$ as $\beta\nearrow\infty$ (see (19)). In our next main result we show that a similar quantization property holds for the radial version of these functionals as well:
Theorem 1.2 (=Theorem 7.3).
Let $\{u_{t}\}\in\mathcal{R}^{1}$. Then
$\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
Given an ample test configuration $\mathcal{T}$, recall that $K_{\mathcal{T}}=K\{u^{\mathcal{T}}_{t}\}$ agrees with the non-Archimedean K energy, which also equals the Futaki invariant of $\mathcal{T}$ if the central fiber of $\mathcal{T}$ is reduced [17, (0.1)]. By the above result, the invariants $K^{\beta}_{\mathcal{T}}$ from (8) converge to this well-known algebro-geometric invariant:
Corollary 1.3.
Let $(\mathcal{X},\mathcal{L})$ be an ample test configuration for $(X,L)$. Then
$\lim_{\beta\to\infty}K^{\beta}_{\mathcal{T}}=K_{\mathcal{T}}.$
We further show in Corollary 8.2 that $K^{\beta}_{\mathcal{T}}$ can be computed algebraically, in terms of log canonical thresholds and intersection numbers associated to $\mathcal{T}$, showing similarities with Berman’s algebraic formula for the slope of the Ding energy along test configurations [5, (3.19)]. At the end of §8 we also establish that, in the Fano case, uniform $K^{\beta}$-stability is equivalent to uniform Ding stability. So in light of Theorem 1.1, our approach also gives a new proof for the uniform YTD conjecture for Fano manifolds. In general, the availability of a threshold-type expressions for $K^{\beta}_{\mathcal{T}}$ suggests promising directions for investigating $K^{\beta}$-stability through algebro-geometric techniques, with the goal of developing novel existence criterions for cscK metrics, in line with recent advances in the Fano setting [60, 53, 54, 16, 1, 3, 90].
@@
Our aforementioned results all rely on a general formula for $Ent^{\beta}\{u_{t}\}$ on bounded geodesic rays, which is a radial version of the identity (4) (Theorem 7.1):
$Ent^{\beta}\{u_{t}\}=\sup_{\{w_{t}\}}\left(L^{\beta}(\{w_{t}\},\{u_{t}\})+\beta(I\{w_{t}\}-I\{u_{t}\})\right),$ (11)
@@
where the sup is over geodesic rays $\{w_{t}\}$ arising from ample test configurations. Here $L^{\beta}(\{w_{t}\},\{u_{t}\})$ is the following generalization of the radial Ding functional:
$L^{\beta}(\{v_{t}\},\{u_{t}\}):=\liminf_{t\to\infty}\left(-\frac{1}{t}\log\int_{X}e^{\beta(v_{t}-u_{t})}\frac{\omega^{n}}{V}\right),$
which can be further expressed as an integrability threshold (see Proposition 7.2).
Non-Archimedean interpretation of $Ent^{\beta}$.
@@
When dealing with non-Archimedean data, it is possible to describe the radial $Ent^{\beta}$ functional using a concise formula. For details on terminology and non-Archimedean geometry we refer to [17] and §8.
@@
Let $u\in\mathcal{H}^{NA}$ be a non-Archimedean metric in the spirit of [17], encoding the data of ample test configurations using divisorial information. As such, to $u$ one can again associate a $C^{1,1}$ geodesic ray $\{u_{t}\}$. As a direct consequence of (11), we obtain the following formula for the radial $\beta$-entropy of $\{u_{t}\}$ (Corollary 9.2):
$Ent^{\beta}\{u_{t}\}=\sup_{w\in\mathcal{H}^{NA}}\left(L^{NA}(\beta(u-w))+\beta(I^{NA}(w)-I^{NA}(u))\right),$ (12)
where $L^{NA}$ and $I^{NA}$ are non-Archimedean functions defined in [20, Section 2].
Since it is purely a non-Archimedean expression, one may define the right hand side of (12) as $Ent^{\beta,NA}(u)$, the non-Archimedean $\beta$-entropy.
It will easily follow from our analysis that the above formula also holds for $u\in\mathcal{H}^{NA}_{\mathcal{M}}$, the non-Archimedean potentials arising from a model (filtration) of C. Li [61, Definition 2.7]. More importantly, using [22, (4.5)] we immediately obtain the inequality $Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(u),\ u\in\mathcal{H}^{NA}_{\mathcal{M}}.$ Now we can use [61, Proposition 6.3] and Theorem 5.2 to conclude that, more generally:
$Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$
@@
Adding the terms coming from the radial Monge-Ampère energy and its twisted version, we immediately obtain that $K^{\beta}\{u_{t}\}\leq K^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$

### v2

@@ Abstract
Given a compact Kähler manifold, to better understand Mabuchi’s $K$ energy we introduce a family of $K^{\beta}$ energies, whose favorable properties are similar to those of the Ding energy from the Fano case. The construction uses Berman’s transcendental quantization, and we show that the slope of the $K^{\beta}$ energies along test configurations can be computed using intersection theory. With these ingredients in place we provide a uniform Yau–Tian–Donaldson correspondence that characterizes the existence of a unique constant scalar curvature Kähler metric using test configurations. Combining our techniques with the non-Archimedean approach to $K$-stability pioneered by Boucksom–Jonsson, we show that the properness of the classical $K$ energy can be tested by checking its slope along a distinguished subclass of Chi Li-type models, called log discrepancy models, thus yielding another $\mathbb{G}$-uniform Yau–Tian–Donaldson correspondence.
@@ 1 The main results
Let $(X,\omega)$ be a compact Kähler manifold. Classical formulations of the Yau–Tian–Donaldson (YTD) conjecture aim to characterize the existence of constant scalar curvature Kähler (cscK) metrics using algebraic invariants associated to test configurations [104, 95, 56, 50, 91, 21, 51, 90]. Our work fits within this general framework. While we do not address any specific version of the YTD conjecture from the literature, we do establish YTD correspondences that characterize existence of cscK metrics using test configurations and novel algebraic invariants.
@@
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X,\mathbb{R})\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
@@
where ${Ent}(u)=\frac{1}{V}\int_{X}\log(\omega^{n}_{u}/\omega^{n})\omega^{n}_{u}$ is the entropy of $\omega^{n}_{u}$ with respect to $\omega^{n}$, whereas ${I},{I}_{\textup{Ric}(\omega)}:\mathcal{H}_{\omega}\to\mathbb{R}$ are the Monge–Ampère (Aubin-Yau) energy and its Ricci-twisted version:
@@
The $K^{\beta}$ energy and transcendental quantization.
@@
Recalling the formula $Ent(u)=\sup_{w\in C^{0}(X)}\left(-\log\frac{1}{V}\int_{X}e^{w}{\omega^{n}}+\frac{1}{V}\int_{X}w\omega_{u}^{n}\right)$ ([7, Proposition 2.10]) and the basic inequality $I(v)-I(u)\leq\frac{1}{V}\int_{X}(v-u)\omega_{u}^{n}$, we immediately see that $Ent^{\beta}(u)\leq Ent(u)$. In Corollary 3.11 we will prove that $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
It may seem surprising at first, but our formulation of the $\beta$-entropy is intimately tied to Berman’s speculations from [5, §3.1], where he suggests to view solutions $u^{\beta}$ to the following Aubin–Yau type Monge–Ampère equation as ‘transcendental quantizations’ of $u\in\mathcal{H}_{\omega}$, with $\beta^{-1}$ serving as the Planck constant:
@@
For our $\beta$-entropy, the following simple but crucial formula is proved in Proposition 3.6:
@@
The expression on the right hand side is reminiscent of Donaldson’s $\tilde{\mathcal{L}}$ function from [57, (12)] in the classical Bergman kernel setting. Because of this analogy, we call $Ent^{\beta}$ the (transcendentally) quantized entropy at level $\beta$. The corresponding quantized $K^{\beta}$ energy functional $K^{\beta}:\mathcal{H}_{\omega}\to\mathbb{R}$—satisfying $K^{\beta}(u)\nearrow K(u)$ as $\beta\nearrow\infty$—is defined as follows:
@@
Due to Chen-Cheng, existence of cscK metrics is equivalent to properness of the K energy [30, 31], which we show to be further equivalent with properness of the ‘quantized’ $K^{\beta}$ functionals for $\beta$ big enough (Theorem 4.4). This is in line with the quantization philosophy extensively explored in the algebraic setting (see e.g. [93, 94, 55, 57, 80, 79, 88, 107]). Our argument relies on estimates reminiscent of the partial $C^{0}$ estimate (Proposition 4.1), originally introduced by Tian [94] in the Fano setting.
In addition, we show that the $K^{\beta}$ energy and its radial version are continuous with respect to the $d_{1}$ and $d_{1}^{c}$ metrics, respectively (Lemma 3.9, Theorem 5.2). The analogous properties are known to hold for the Ding energy ([43, Proposition 5.19], [45, Theorem 1.4] [25, Proposition 3.15]) but fail for the $K$ energy. Thus, while the properness profile of $K^{\beta}$ agrees with that of the $K$ energy, its crucial continuity profile is in line with the much better behaving Ding energy from the Fano case.
Since, in the Fano case, the favorable properties of the Ding energy have driven important advances in $K$-stability, moduli theory, $\delta$-invariants, and other areas [13, 32, 96, 68, 58, 59, 15, 102, 73, 107, 46], we are motivated to make extensive use of $K^{\beta}$ energies in this work.
A YTD correspondence using $K^{\beta}$.
Given an ample line bundle $L\to X$ with hermitian metric $h$ whose curvature equals our Kähler metric $\omega$, to test K-stability, one considers test configurations. Historically this notion is due to Ding–Tian and Donaldson, and involves a one parameter degeneration of the structure $(X,L)$ [54, 95, 56]. We give the precise definition of ample test configurations in §2. One may think of these as certain finitely generated $\mathbb{Z}$-filtrations of the section ring $R(X,L^{k})$ for some $k\in\mathbb{N}$. This perspective was first considered in [98, 91] and it was shown in [21, Proposition 2.15] that it is equivalent with the more classical notion introduced in [95, 56].
To an ample test configuration $\mathcal{T}$, due to work initiated by Phong-Sturm, it is possible to associate a $C^{1,1}$ Mabuchi geodesic ray $0,\infty)\ni t\to u^{\mathcal{T}}_{t}\in\mathcal{H}_{\omega}^{1,1} [$82, 83, 86, 4, 34].
@@
Given that $\mathcal{T}$ is algebraic, it is natural to seek for an interpretation of the associated invariants purely in algebraic terms. For instance, $J_{\mathcal{T}}$ admits such a description via certain intersection numbers attached to $\mathcal{T}$ [81, 21]. For $K^{\beta}_{\mathcal{T}}$, an explicit intersection-theoretic formula will be provided in Theorem 1.2 below. In §7 we also provide a threshold type expression for $K^{\beta}_{\mathcal{T}}$, that may open the door to new valuative criterions for K-stability, generalizing those in the Fano case [58, 59, 15, 1, 2, 101].
@@
To explain the proof of Theorem 1.1, let us recall a few concepts. By $(\mathcal{E}^{1},d_{1})$ we denote the $d_{1}$ metric completion of $\mathcal{H}_{\omega}$, explored in [36], with the strongly topology of the space $\mathcal{E}^{1}$ already investigated in [7]. The space $(\mathcal{E}^{1},d_{1})$ is a geodesic metric space [36, Theorem 2]. Our functionals $K$, $K^{\beta}$, and $J$ extend naturally to $\mathcal{E}^{1}$, preserving their $d_{1}$-(semi)continuity properties [6, 10]. Moreover, the extended $K^{\beta}$ continues to satisfy the formulas (4) and (7).
By $\mathcal{R}^{1}$ we denote the space of geodesic rays $t\mapsto u_{t}$ inside $\mathcal{E}^{1}$, that emanate from the background potential $0\in\mathcal{H}_{\omega}$ (notation: $\{u_{t}\}\in\mathcal{R}^{1}$). Due to Buseman convexity of $d_{1}$ ([10, Proposition 5.1]) one can introduce the chordal distance for $\{u_{t}\},\{v_{t}\}\in\mathcal{R}^{1}$ [41, 31]:
@@
Thus defined, $(\mathcal{R}^{1},d_{1}^{c})$ is a complete geodesic metric space [41, Theorem 1.4].
An important subspace of $(\mathcal{R}^{1},d_{1}^{c})$ is $\mathcal{R}^{1}_{\mathcal{I}}$, defined as the $d_{1}^{c}$-closure of rays $\{u_{t}^{\mathcal{T}}\}$ induced by test configurations $\mathcal{T}$. In [9] the authors refer to such rays simply as maximal. This space admits a natural identification with $\mathcal{E}^{1,NA}$ (see [44, Theorem 1.2] and the discussion preceding it). For more detailed treatments of this topic, we refer to [41, 9, 44, 38].
Given $\{u_{t}\}\in\mathcal{R}^{1}$ it is convenient to introduce the radial version of our functionals:
@@
In a recent exciting development, Chi Li showed that rays $\{u_{t}\}\in\mathcal{R}^{1}$ satisfying $K\{u_{t}\}<\infty$ must lie in $\mathcal{R}^{1}_{\mathcal{I}}$ [70, Theorem 1.2].
The proof of Theorem 1.1 relies on four central components. From the existing literature, we will need that existence cscK metrics is equivalent with properness/geodesic stability of the K energy [30, 31], together with the aforementioned result of Chi Li [70]. The remaining two components are facts we establish in this work: that the properness of the K energy is equivalent to the properness of the $K^{\beta}$ energy for sufficiently large $\beta$ (Theorem 4.4); and that the radial $K^{\beta}$ energy is $d_{1}^{c}$-continuous (Theorem 5.2).
Once the key ingredients are in place, the argument unfolds naturally. We present it here, as this simple proof serves as the skeleton of the paper (cf. [9, 58] for related ideas in the Fano case involving the Ding energy).
@@
Assume that (i) holds. By [11, Theorem 1.5] the K energy is proper, i.e., $K(u)\geq\gamma J(u)-C,\ u\in\mathcal{E}^{1}$ for some $\gamma,C$. By Theorem 4.4 we obtain that $K^{\beta}$ is proper too, for $\beta>0$ big enough. By our definition of $K^{\beta}_{\mathcal{T}}$ and $J_{\mathcal{T}}$, (ii) immediately follows.
Now we assume that (ii) holds. By [70, Theorem 1.2] any ray with bounded K energy slope is in $\mathcal{R}^{1}_{\mathcal{I}}$, i.e., it is approximable by test configurations. Thus, by [31] (cf. [41, Theorem 1.9]), it suffices to show existence of $\gamma>0$ such that $K\{u_{t}\}\geq\gamma J\{u_{t}\}$ for $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}.$ Let $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}$ and $\{u^{\mathcal{T}_{j}}_{t}\}$ be $d_{1}^{c}$-approximating rays of ample test configurations $\mathcal{T}_{j}$.
@@
The functional $J\{\cdot\}$ is known to be $d_{1}^{c}$-continuous (as follows from [36, Lemma 4.15] and [9, Lemma A.2]). $K^{\beta}\{\cdot\}$ also has this property by Theorem 5.2. Thus, using $K\geq K^{\beta}$ (see (33)), we can let $j\to\infty$ in the above inequality to obtain $K\{u_{t}\}\geq K^{\beta}\{u_{t}\}\geq\gamma J\{u_{t}\}.$
Algebraic interpretation of $K^{\beta}_{\mathcal{T}}$.
As Theorem 1.1 introduces a new stability condition in the Yau–Tian–Donaldson framework, the second goal of this paper is to better understand the meaning of $K^{\beta}$-stability and connect it with more classical notions of stability.
To this end, we shall work with the slightly more general class of semi-ample test configurations, which are just pull-backs of ample ones (see §2) and are birationally more convenient to work with. Given such test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$, after possibly composing with blowups, we can further assume that $\mathcal{X}$ is smooth and dominates $X\times{\mathbb{P}}^{1}$ via a map $\pi$.
Then $K_{\mathcal{T}}:=K\{u^{\mathcal{T}}_{t}\}$ agrees with the non-Archimedean K energy [22], [70, (8)]:
$K_{\mathcal{T}}=V^{-1}\mathcal{L}^{n}\cdot K^{\log}_{\mathcal{X}/X\times{\mathbb{P}}^{1}}+\frac{\bar{S}\mathcal{L}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\mathcal{L}^{n}}{V},$ (12)
which equals the Futaki invariant of $\mathcal{T}$ if the central fiber of $\mathcal{X}$ is reduced [21, (0.1)]. Here
$K^{\log}_{\mathcal{X}/X\times{\mathbb{P}}^{1}}:=K_{\mathcal{X}}-\pi^{*}K_{X\times{\mathbb{P}}^{1}}+\mathcal{X}_{0,\mathrm{red}}-\mathcal{X}_{0}$
is the relative log canonical divisor (see [21, §4.4]).
Next we show that $K^{\beta}_{\mathcal{T}}$ also admits a purely algebraic expression. The main difficulty is that the explicit expression for $K^{\beta}$ (see (7) and (8)) involves solving the transcendental Aubin–Yau type equation (5). This will be overcome using the next key concept.
Given a smooth test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$ dominating $X\times{\mathbb{P}}^{1}$, we call $\mathcal{T}_{\beta}:=(\mathcal{X},\mathcal{L}_{\beta})$ the log discrepancy model of $\mathcal{T}$ at level $\beta$, where
$\mathcal{L}_{\beta}:=\mathcal{L}+\frac{1}{\beta}K^{\log}_{\mathcal{X}/X\times\mathbb{P}^{1}},\ \beta\in\mathbb{Q}_{>1}.$ (13)
The $\mathbb{Q}$-line bundle $\mathcal{L}_{\beta}$ thus defined is typically not relatively semi-ample, making $\mathcal{T}_{\beta}=(\mathcal{X},\mathcal{L}_{\beta})$ a model of $(X,L)$ in the sense of Chi Li [70, Definition 2.1] (see §2).
Theorem 1.2 (Theorem 9.5).
For a test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$ as above, and $\beta\in\mathbb{Q}_{>1}$, we have the following:
$K^{\beta}_{\mathcal{T}}=\frac{\beta(\langle\mathcal{L}_{\beta}^{n+1}\rangle-\mathcal{L}^{n+1})}{(n+1)V}+\frac{\bar{S}\mathcal{L}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\mathcal{L}^{n}}{V}.$ (14)
Moreover, $K^{\beta}_{\mathcal{T}}\nearrow K_{\mathcal{T}}$ as $\beta\nearrow\infty$.
The term $\langle\mathcal{L}_{\beta}^{n+1}\rangle$ appearing in (14) denotes the volume of the (relatively) big $\mathbb{Q}$-line bundle $\mathcal{L}_{\beta}$, which can be expressed using movable intersection products (see (70) and [18, 71]).
@@
The proof of Theorem 1.2 relies on a slope formula for $Ent^{\beta}\{u_{t}\}:=\liminf_{t\to\infty}Ent^{\beta}(u_{t})/t$, for any bounded geodesic ray $\{u_{t}\}$. It is a radial version of the identity (4) (Theorem 6.1):
@@
where the sup is over geodesic rays $\{v_{t}\}$ arising from test configurations, and $L^{\beta}(\{v_{t}\},\{u_{t}\})$ denotes the following radial Ding type functional:
$L^{\beta}(\{v_{t}\},\{u_{t}\}):=\liminf_{t\to\infty}\left(-\frac{1}{t}\log\int_{X}e^{\beta(v_{t}-u_{t})}\frac{\omega^{n}}{V}\right).$
To prove (14), we will show in §9 that, when the ray $\{u_{t}\}$ is induced by a test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$, the sup in (15) is actually attained by the geodesic ray $\{v_{t}\}$ associated with the log discrepancy model $\mathcal{T}_{\beta}=(\mathcal{X},\mathcal{L}_{\beta})$.
Interactions with the non-Archimedean point of view.
In the last part of our paper we point out the connections between our $K^{\beta}$-stability and the non-Archimedean approach to K-stability pioneered by Boucksom–Jonsson and collaborators [9, 21, 23, 26, 25, 27].
@@
For details on terminology and non-Archimedean geometry we refer to [21] and §2. When dealing with non-Archimedean data, it is possible to describe the radial $Ent^{\beta}$ functional using a concise formula.
@@
Let $u\in\mathcal{H}^{NA}$ be a non-Archimedean metric, as defined in [21], encoding the data of test configurations using divisorial information. As such, to $u$ one can again associate a $C^{1,1}$ geodesic ray $\{u_{t}\}$. As a consequence of (15), we obtain the following formula for the radial $\beta$-entropy of $\{u_{t}\}$ (Corollary 8.2):
$Ent^{\beta}\{u_{t}\}=\sup_{w\in\mathcal{H}^{NA}}\left(L^{NA}(\beta(u-w))+\beta(I^{NA}(w)-I^{NA}(u))\right).$ (16)
where $L^{NA}$ and $I^{NA}$ are non-Archimedean functions defined in [23, Section 2]. As we will show in §8, due to basic non-Archimedean formalism, the above further implies that
$Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(u),\ \ u\in\mathcal{E}^{1,NA},$
@@
where $Ent^{NA}(u)$ denotes the non-Archimedean entropy introduced in [21, 23]. Adding the terms coming from the radial Monge–Ampère energy and its twisted version, we immediately obtain that $K^{\beta}\{u_{t}\}\leq K^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$